OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (anonymous):

\[\left| 2 x+6 \right|-4=20\] \[\left| 2 x+6 \right|=20+4=24\] \[2 x+6=\pm 24\] x=?

OpenStudy (anonymous):

the answer to the question is C :)

OpenStudy (anonymous):

are you sure I have to get this one right

OpenStudy (anonymous):

yeah I checked on my test. I took the same one and I had this question right

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

2 x+6=24 2 x=24-6=18 x=18/2=9 2x+6=-24 2x=-24-6=-30 x=-30/2=-15